# EffoVPR: Effective Foundation Model Utilization for Visual Place Recognition

### v1

@@ 3.1 Training strategy
Our training strategy employs a classification loss applied on the output $[CLS]$ token. Following EigenPlaces [11], we partition the map into cells measuring $15\times 15$ meters each, where cells define classes. Each class contains images capturing the same location from various viewpoints, enhancing the model’s ability to accurately recognize locations despite substantial changes in the captured view angle. To enforce class discrimination while ensuring viewpoint invariance, we utilize the Large Margin Cosine Loss (CosFace)[8, 45].
We initialize our ViT[17] backbone with pre-trained DINOv2 weights [37, 16]. To retain the rich visual representations learned during pre-training, while adapting the model for the VPR task, we fine-tune only the final layers of our backbone. Note that due to the inherent design of the ViT architecture, which incorporates a self-attention mechanism (Eq. 1), the global feature is implicitly trained to aggregate local features without the need of additional specialized components. For learning more compact global feature we simply add a linear layer on top of the output [CLS] token.
@@ 3.2 Inference strategy
Re-ranking stage: We start by extracting patch local features from each candidate among the top-k retrieved images in the previous global stage, and re-rank the candidates based on these features. Local features are derived from the intermediate layer $l$ of ViT, utilizing the self-attention matrices, thus their computation is integrated into the process of computing global features and does not necessitate recalculation. Aligned with the findings in [3], where $V_{l}$ showed to have a higher instance-level characteristics, we find $V_{l}$ matrix mentioned in Eq. 1 as the most propitious facet for the task at hand (see Tab. S3 in Appx.). Then we leverage the model’s internal prioritization to identify discriminative features and extract the attention map: $S:=\text{Softmax}(Q_{l}\cdot k_{cls})\in\mathbb{R}^{p}$, which represents the attention of each Value feature with the global $[CLS]$ image representation key-token $k_{cls}$. We therefore select a subset $\mathcal{V}$ of the values $\mathcal{V}\subseteq\{v_{1},...,v_{p}\}:=V_{l}$ as the image’s local features, based on the score $S$: $\mathcal{V}:=\{v_{i}~{}|~{}S_{i}>T_{1}\}$ for a predefined threshold $T_{1}$. We use a rather low threshold that filters out patches that are “less significant” as reflected from the layer’s attention. Note that the number of selected local features might differ between images.
Next, given the query’s local features $\mathcal{V}$ and a candidate’s $\mathcal{V}^{\prime}$, we calculate the pairs of mutual nearest neighbors (MNN) between $\mathcal{V}$ and $\mathcal{V}^{\prime}$ by the number of feature pairs that are the nearest neighbor of each other. Our observation revealed that applying a threshold to the MNN scores enhances the model’s resilience to clutter and directs its attention to the pertinent key-points for accurate matching. We therefore count only the pairs with cos-similarity higher than a predefined threshold $T_{2}$ (for ablation on thresholds see Tab. S3 in Appx.). We formulate that process in Equation 2:
$\text{MNN}(\mathcal{V},\mathcal{V}^{\prime}):=\{(v_{i},v^{\prime}_{j})\in\mathcal{V}\times\mathcal{V}^{\prime}~{}|~{}v_{i}:=NN(V,v^{\prime}_{j})~{}\text{and}~{}v^{\prime}_{j}:=NN(v_{i},V^{\prime}),~{}v_{i}^{T}v^{\prime}_{j}>T_{2}\}$ (2)
@@ 4 Evaluation
In this section, we compare our approach with several SoTA VPR methods following the common VPR Benchmarks[10]. We propose, a single-stage (EffoVPR-G) and two-stage, that includes a re-ranker (EffoVPR-R) approaches with backbone trained on the publicly available SF-XL [8] dataset containing streetview of San Francisco. For more implementation details see Appx. We then test EffoVPR on a large number of diverse datasets (20), including e.g. Pitts30k [4], Tokyo24/7[43], MSLS-val/challenge[47] Nordland[41] and more, exhibiting a wide variety of conditions, including different cities, day/night images, and seasonal changes. Note that, the MSLS Challenge[47] is a hold-out set whose labels are not released, but researchers submit the predictions to the challenge server to get the performance. More details on the benchmarks can be found in the Appendix.
Datasets with gallery made from street-view images and with largest viewpoint variance, include Tokyo 24/7[43] and SF-XL[45, 6], where the query images are collected from a smartphone, usually from sidewalks. Most datasets are from urban footage, with the main exception being Nordland[41], which is a collection of photos taken across different seasons with a camera mounted on a train. Some datasets present various degrees of day-to-night changes, namely MSLS [47], Tokyo 24/7 [43], SF-XL [8] SVOX-Night [12]. AmsterTime [48] contains grayscale historical queries and modern-time RGB gallery images, making it the only dataset with large-scale time variations of multiple decades.
We follow common evaluation metric used in previous works e.g. [10, 11, 49, 34, 33, 4] and use 25 meters radius as the threshold for correct localization and report Recall@K metrics for K=1,5,10. For Nordland we evaluate ±$\pm 10$ frames as the common evaluation protocol used in [10, 21]. For a more comprehensive description of all 20 datasets and implementation details see Appx.
@@ 4.1 Zero-shot performance
We present the performance of our zero-shot method (EffoVPR-ZS) in Table 1 compared to two zero-shot alternatives, the recently published AnyLoc, and DINOv2 global feature (using the output [CLS] token), without finetuning, where EffoVPR-ZS re-ranks its top-100 global retrieved candidates. The results show that our method significantly improves over the baseline and is superior to AnyLoc. Note the significant gap for more challenging scenarios of Tokyo24/7 and Nordland exhibiting day vs. night and seasonal variations. AnyLoc tends to fail in these challegning scenarios as its VLAD aggregation learned (in unsupervised manner) on the gallery can not generalize well to challenging out-of-distribution queries. In Figure 3(a) we compare our zero-shot approach with several methods that have used VPR datasets for training. Although EffoVPR-ZS was not trained on VPR task, it still achieves comparable results to the trained methods on three popular datasets. This success can be attributed to the robust features in DINOv2, specifically those selected from the $\mathcal{V}$ facet, combined with our mutual-NN matching and scoring. Figure 3(b) demonstrates this by showcasing a scenario where the original attention, mistakenly focusing on an advertisement placed in front of a building. However, our method successfully identifies relevant key-points on the building itself, enabling correct image matching (even though there is a different ad on the gallery image).
@@ 4.2 Comparison with State-of-The-Art
In this section, we compare our single stage (EffoVPR-G) and two-stage methods (EffoVPR-R) with previous state-of-the-art including the recent works of [34, 22, 33]. SelaVPR and R2Former were trained on a combination of Pitts30k and MSLS while CricaVPR and SALAD were trained on GSV-Cities, and Cosplace and EigenPlaces on SF-XL (similar to ours). We show in Table 2 the global retrieval results (without re-ranking) with Recall@1 on five different benchmarks. EffoVPR-G achieves SoTA performance on three out of five datasets, while being ranked second on the other two. This highlights the effectiveness of the single global representation learned by our method. Notably, it achieves $+2.9\%$ on Tokyo24/7, $+2.8\%$ on the challenging Nordland dataset that exhibits extreme seasonal changes, and $+3.2\%$ on the hold-out MSLS-challenge dataset. We further demonstrate the performance of our global feature learning with reduced dimensions of 256D and even 128D, significantly decreasing the memory footprint and enabling efficient searches within a considerably larger gallery. The findings indicate only a marginal degradation in performance with lower-dimensional features, while achieving parity with the SALAD on Tokyo24/7 using 128D, compared to 8,448D, feature size (a 66-fold reduction in dimensionality). Figure 1 illustrates this quality on Tokyo24/7 and the hold-out MSLS-challenge, showing our top performing results even with 128D feature size. Results on more datasets can be found in the Appendix. The strong impact of our finetuning process including the last five layers, is visualized in Figure 4.
@@
CosPlace [8] 512 90.5 81.9 82.8 61.4 66.5
MixVPR [2] 4096 91.5 86.7 88.2 64.0 58.4
R2Former [49] 256 76.3 45.7 79.3 56.2 50.9
EigenPlaces [11] 2048 92.5 93.0 89.1 67.4 71.2
SelaVPR [34] 1024 90.2 81.9 87.7 69.6 72.3
CricaVPR [33] 4096 94.9* 93.0 90.0 69.0 90.7
SALAD [22] 8448 92.4 94.6 92.2 75.0 76.0
@@
Next, we showcase the comprehensive performance of our two-stage approach (EffoVPR-R) in Table 3. EffoVPR-R achieves top performance across all datasets, taking the second place only on Pitts30K-R@1, with very close result. Note that CricaVPR reports using Pitts30k as a validation set, which may have contributed to the improved results on this dataset. However, EffoVPR-R demonstrates notable improvements, particularly evident in Tokyo24/7, where it achieves a remarkable increase in R@1 from 94.6% to 98.7% and in MSLS-challenge (from to 75.0% to 79.0%). These results underscore the generalization capability of our approach, demonstrating its resilience in handling significant variations between query and gallery images, such as viewpoint discrepancies (as seen in Pitts30k) and changes in illumination (in Tokyo24/7), across a diverse range of locations. Following the common practice, we report EffoVPR-R re-ranking performance over the top-100 candidates retrieved in the first-stage ($K=100$), however we achieve SoTA R@1 results even with a low number of candidates (from $K=5$ onwards, see Tab. S4 in Appx). Note that EffoVPR matching method is highly expedient, with an average processing time of just 1 millisecond per match.
@@
NetVLAD [4] 81.9 91.2 93.7 60.6 68.9 74.6 53.1 66.5 71.1 35.1 47.4 51.7
SFRS [20] 89.4 94.7 95.9 81.0 88.3 92.4 69.2 80.3 83.1 41.6 52.0 56.3
Patch-NetVLAD [21] 88.7 94.5 95.9 86.0 88.6 90.5 79.5 86.2 87.7 48.1 57.6 60.5
CosPlace [8] 88.4 94.5 95.7 81.9 90.2 92.7 82.8 89.7 92.0 61.4 72.0 76.6
MixVPR [2] 91.5 95.5 96.4 86.7 92.1 94.0 88.2 93.1 94.3 64.0 75.9 80.6
R2Former [49] 91.1 95.2 96.3 88.6 91.4 91.7 89.7 95.0 96.2 73.0 85.9 88.8
EigenPlaces [11] 92.5 96.8 97.6 93.0 96.2 97.5 89.1 93.8 95.0 67.4 77.1 81.7
SelaVPR [34] 92.8 96.8 97.7 94.0 96.8 97.5 90.8 96.4 97.2 73.5 87.5 90.6
CricaVPR [33] 94.9 97.3 98.2 93.0 97.5 98.1 90.0 95.4 96.4 69.0 82.1 85.7
SALAD [22] 92.4 96.3 97.4 94.6 97.5 97.8 92.2 96.2 97.0 75.0 88.8 91.3
@@
Finally, we demonstrate the efficacy of our approach in the most challenging VPR benchmark scenarios by conducting experiments on six demanding datasets: Nordland [41], which includes extensive seasonal changes; AmsterTime [48], spanning over an extended time period; SF-Occlusion[6], that features queries with significant field-of-view obstructions; SF-Night[6], with severe illumination changes; and SVOX [12], with extreme weather and illumination variations. The results, detailed in Table 4, underscore the significant superiority of our method over previous approaches across these datasets. EffoVPR-R shows improvements of +4.3%, +0.8%, +7.9%, and +15%, +2% on Nordland, AmsterTime, SF-Occlusion, SF-Night, and SVOX-Night respectively and comparable results on SVOX-Rain. This demonstrates the high versatility of our model, which can handle extreme variations even when trained without seasonal or day-to-night changes. Figure 1 shows an examples of this case. We attribute this robustness primarily to the combination of our training method and specific re-ranking strategy over the DINOv2 model. We conduct an extensive ablation study on various hyperparameters and aspects of our approach in the Appendix.
@@
EigenPlaces [11] 71.2 48.9 32.9 23.6 58.9 90.0
SelaVPR [34] 85.2 55.2 35.5 38.4 89.4 94.7
CircaVPR [33] 90.7 64.7 42.1 35.4 85.1 95.0
SALAD [22] 76.0 58.8 51.3 46.6 95.4 98.5
@@
5 Summary and Limitations
In this paper, we introduced a single and two-stage approach for VPR, that effectively leverages a foundation model. Our method utilizes existing internal self-attention and pooling mechanisms to propose a new approach that achieves high performance even in a zero-shot setting.
We observed that despite the success of our zero-shot approach, it does not grasps the relevance of certain objects in the scene, for localization. This limitation is highlighted in Figure 5, where the iconic and visually identical cable car in SF causes distraction. Fine-tuning the model resolves this problem by shifting the attention from transient foreground objects to static VPR-relevant cues, as seen in Figure 4. Nevertheless, while our second-stage matching approach proves highly effective, we forgo geometric verification for sake of speed. Integrating such approach, in the future may offer further refinement and performance enhancement.
The experimental results demonstrated that our trained model outperforms previous SoTA often by a large margin, particularly in demanding scenarios that exhibit strong appearance change. Having compact features, our method provides a promising way to address the VPR task in real-world, large-scale applications.

### v2

@@ 3.1 Training strategy
Our training strategy employs a classification loss applied on the output $[CLS]$ token. Following EigenPlaces (Berton et al., 2023), we partition the map into cells measuring $15\times 15$ meters each, where cells define classes. Each class contains images capturing the same location from various viewpoints, enhancing the model’s ability to accurately recognize locations despite substantial changes in the captured view angle. To enforce class discrimination while ensuring viewpoint invariance, we utilize Large Margin Cosine Loss (CosFace) (Berton et al., 2022a; Wang et al., 2018).
We initialize our ViT (Dosovitskiy et al., 2021) backbone with pre-trained DINOv2 weights (Oquab et al., 2023; Darcet et al., 2023). To retain the rich visual representations learned during pre-training, while adapting the model for the VPR task, we fine-tune only the final layers of our backbone. Note that due to the inherent design of the ViT architecture, which incorporates a self-attention mechanism (Eq. 1), the global feature is implicitly trained to aggregate local features without the need of additional specialized components. For learning more compact global feature we simply add a linear layer on top of the output [CLS] token.
@@ 3.2 Inference strategy
Re-ranking stage: We start by extracting patch local features from each candidate among the top-k retrieved images in the previous global stage, and re-rank the candidates based on these features. Local features are derived from the intermediate layer $l$ of ViT, utilizing the self-attention matrices, thus their computation is integrated into the process of computing global features and does not necessitate recalculation. For keypoints descriptors, we find that $V_{l}$ matrix in Eq. 1 most effectively captures the ”instance” properties of the keypoints, making it the most promising facet for the task at hand (see analysis in Tab. S3 in Appendix). Then, for keypoint selection, we leverage the model’s internal prioritization, to identify discriminative features and extract the attention map: $S:=\text{Softmax}(Q_{l}\cdot k_{cls})\in\mathbb{R}^{p}$, which represents the attention of each Value feature with the global $[CLS]$ image representation key-token $k_{cls}$. We therefore select a subset $\mathcal{V}$ of the values $\mathcal{V}\subseteq\{v_{1},...,v_{p}\}:=V_{l}$ as the image’s local features, based on the score $S$: $\mathcal{V}:=\{v_{i}~{}|~{}S_{i}>T_{1}\}$ for a predefined threshold $T_{1}$. Note that the number of selected local features may vary between images.
Next, given the query’s local features $\mathcal{V}$ and a candidate’s $\mathcal{V}^{\prime}$, we calculate the pairs of mutual nearest neighbors (MNN) between $\mathcal{V}$ and $\mathcal{V}^{\prime}$ by the number of feature pairs that are the nearest neighbor of each other. Our observation revealed that applying a threshold to the MNN scores enhances the model’s resilience to clutter and directs its attention to the pertinent key-points for accurate matching. We therefore count only the pairs with cos-similarity higher than a predefined threshold $T_{2}$. We formulate that process in Equation 2:
$\text{MNN}(\mathcal{V},\mathcal{V}^{\prime}):=\{(v_{i},v^{\prime}_{j})\in\mathcal{V}\times\mathcal{V}^{\prime}~{}|~{}v_{i}:=NN(V,v^{\prime}_{j})~{}\text{and}~{}v^{\prime}_{j}:=NN(V^{\prime},v_{i}),~{}v_{i}^{T}v^{\prime}_{j}>T_{2}\}$ (2)
@@
Our design parameters include the two key values of $T_{1},T_{2}$. Threshold $T_{1}$ is used to select strong keypoints for image matching, typically focusing on more significant areas. The threshold $T_{2}$ facilitates the selection of strong matching pairs, filtering out weak correspondences that could result in false image matches. For a detailed ablation study on these thresholds refer to Appendix B.
@@ 4 Evaluation
In this section, we compare our approach with several SoTA VPR methods following the common VPR Benchmarks (Berton et al., 2022c). We propose, a single-stage (EffoVPR-G) and two-stage, that includes a re-ranker (EffoVPR-R) approaches with backbone trained on the publicly available SF-XL (Berton et al., 2022a) dataset containing streetview of San Francisco. For more implementation details see Appendix. We then test EffoVPR on a large number of diverse datasets (20), including e.g. Pitts30k (Arandjelovic et al., 2016), Tokyo24/7 (Torii et al., 2015), MSLS-val/challenge (Warburg et al., 2020) Nordland (Sünderhauf et al., 2013) and more, exhibiting a wide variety of conditions, including different cities, day/night images, and seasonal changes. Note that MSLS Challenge (Warburg et al., 2020) is a hold-out set whose labels are not released, but researchers submit the predictions to the challenge server. Further details on the benchmarks are in the Appendix.
Datasets with gallery made from street-view images and with largest viewpoint variance, include Tokyo 24/7 (Torii et al., 2015) and SF-XL (Wang et al., 2018; Barbarani et al., 2023), where the query images are collected from a smartphone, usually from sidewalks. Most datasets are from urban footage, with the main exception being Nordland (Sünderhauf et al., 2013), which is a collection of photos taken across different seasons with a camera mounted on a train. Some datasets present various degrees of day-to-night changes, namely MSLS (Warburg et al., 2020), Tokyo 24/7 (Torii et al., 2015), SF-XL (Berton et al., 2022a) SVOX-Night (Berton et al., 2021). AmsterTime (Yildiz et al., 2022) contains grayscale historical queries and modern-time RGB gallery images, making it the only dataset with large-scale time variations of multiple decades.
We follow common evaluation metric used in previous works e.g. (Berton et al., 2022c; 2023; Zhu et al., 2023; Lu et al., 2024b; a; Arandjelovic et al., 2016) and use 25 meters radius as the threshold for correct localization and report Recall@K metrics for K=1,5,10. For Nordland we evaluate ±$\pm 10$ frames as the common evaluation protocol used in (Berton et al., 2022c; Hausler et al., 2021). For a more comprehensive description of all 20 datasets and implementation details see Appendix.
@@ 4.1 Zero-shot performance
We present the performance of our zero-shot method (EffoVPR-ZS) in Table 1 compared to two zero-shot alternatives, the recently published AnyLoc, and DINOv2 global feature (using the output [CLS] token), without finetuning, where EffoVPR-ZS re-ranks its top-100 global retrieved candidates. The results show that our method significantly improves over the baseline and is superior to AnyLoc. Note the significant gap for more challenging scenarios of Tokyo24/7 and Nordland exhibiting day vs. night and seasonal variations. AnyLoc tends to fail in these challegning scenarios as its VLAD aggregation learned (in unsupervised manner) on the gallery can not generalize well to challenging out-of-distribution queries. In Figure 3(a) we compare our zero-shot approach with several methods that have used VPR datasets for training. Although EffoVPR-ZS was not trained on VPR task, it still achieves comparable results to the trained methods on three popular datasets. This success can be attributed to the robust features in DINOv2, specifically those selected from the $\mathcal{V}$ facet, combined with our mutual-NN matching and scoring. Figure 3(b) demonstrates this by showcasing a scenario where the original attention mistakenly focusing on an advertisement placed in front of a building. However, our method successfully identifies relevant key-points on the building itself, enabling correct image matching (even though a different advertisement is displayed on the gallery image).
@@ 4.2 Comparison with State-of-The-Art
In this section, we compare our single stage (EffoVPR-G) and two-stage methods (EffoVPR-R) with previous state-of-the-art including the recent works of (Lu et al., 2024a; Izquierdo & Civera, 2024; Lu et al., 2024b; Ali-bey et al., 2024). SelaVPR and R2Former were trained on a combination of Pitts30k and MSLS while CricaVPR, SALAD and BoQ were trained on GSV-Cities, and Cosplace and EigenPlaces on SF-XL (similar to ours). We show in Table 2 the global retrieval results (without re-ranking) with Recall@1 on five different benchmarks. EffoVPR-G achieves SoTA performance on three out of five datasets, while being ranked second on the other two. This highlights the effectiveness of the single global representation learned by our method. Notably, it achieves $+2.9\%$ on Tokyo24/7, $+2.8\%$ on the challenging Nordland dataset that exhibits extreme seasonal changes, and $+3.2\%$ on the hold-out MSLS-challenge dataset. We further demonstrate the performance of our global feature learning with reduced dimensions of 256D and even 128D, significantly decreasing the memory footprint and enabling efficient searches within a considerably larger gallery. The findings indicate only a marginal degradation in performance with lower-dimensional features, while achieving parity with the SALAD on Tokyo24/7 using 128D, compared to 8,448D, feature size (a 66-fold reduction in dimensionality). Figure 1 illustrates this quality on Tokyo24/7 and the hold-out MSLS-challenge, showing our top performing results even with 128D feature size. Results on more datasets can be found in the Appendix. The strong impact of our finetuning process including the last five layers, is visualized in Figure 4.
@@
CosPlace 512 88.4 81.9 82.8 61.4 66.5
MixVPR 4096 91.5 86.7 88.2 64.0 58.4
R2Former 256 76.3 45.7 79.3 56.2 50.9
EigenPlaces 2048 92.5 93.0 89.1 67.4 71.2
SelaVPR 1024 90.2 81.9 87.7 69.6 72.3
CricaVPR 4096 94.9* 93.0 90.0 69.0 90.7
SALAD 8448 92.4 94.6 92.2 75.0 76.0
BoQ 16384 92.5 91.1 90.7 73.3 83.4
@@
Next, we showcase the comprehensive performance of our two-stage approach (EffoVPR-R) in Table 3. EffoVPR-R achieves top performance across all datasets, taking the second place only on Pitts30K-R@1, with very close result. Note that CricaVPR reports using Pitts30k as a validation set, which may have contributed to the improved results on this dataset. However, EffoVPR-R demonstrates notable improvements, particularly evident in Tokyo24/7, where it achieves a remarkable increase in R@1 from 94.6% to 98.7% and in MSLS-challenge (from to 75.0% to 79.0%). These results underscore the generalization capability of our approach, demonstrating its resilience in handling significant variations between query and gallery images, such as viewpoint discrepancies (as seen in Pitts30k) and changes in illumination (in Tokyo24/7), across a diverse range of locations. Following the common practice, we report EffoVPR-R re-ranking performance over top-100 candidates retrieved in the first-stage ($K=100$), however we achieve SoTA R@1 results even with a low number of candidates (from $K=5$ onwards, see Tab. S4 in Appx). Note that EffoVPR matching method is highly expedient, with an average processing time of just 1 millisecond per match.
@@
NetVLAD 81.9 91.2 93.7 60.6 68.9 74.6 53.1 66.5 71.1 35.1 47.4 51.7
SFRS 89.4 94.7 95.9 81.0 88.3 92.4 69.2 80.3 83.1 41.6 52.0 56.3
Patch-NetVLAD 88.7 94.5 95.9 86.0 88.6 90.5 79.5 86.2 87.7 48.1 57.6 60.5
CosPlace 88.4 94.5 95.7 81.9 90.2 92.7 82.8 89.7 92.0 61.4 72.0 76.6
MixVPR 91.5 95.5 96.4 86.7 92.1 94.0 88.2 93.1 94.3 64.0 75.9 80.6
R2Former 91.1 95.2 96.3 88.6 91.4 91.7 89.7 95.0 96.2 73.0 85.9 88.8
EigenPlaces 92.5 96.8 97.6 93.0 96.2 97.5 89.1 93.8 95.0 67.4 77.1 81.7
SelaVPR 92.8 96.8 97.7 94.0 96.8 97.5 90.8 96.4 97.2 73.5 87.5 90.6
CricaVPR 94.9 97.3 98.2 93.0 97.5 98.1 90.0 95.4 96.4 69.0 82.1 85.7
SALAD 92.4 96.3 97.4 94.6 97.5 97.8 92.2 96.2 97.0 75.0 88.8 91.3
BoQ 92.5 96.4 97.4 91.1 95.9 96.5 90.7 94.7 95.8 73.3 82.9 86.2
@@
Finally, we demonstrate the efficacy of our approach in the most challenging VPR benchmark scenarios by conducting experiments on six demanding datasets: Nordland (Sünderhauf et al., 2013), which includes extensive seasonal changes; AmsterTime (Yildiz et al., 2022), spanning over an extended time period; SF-Occlusion (Barbarani et al., 2023), that features queries with significant field-of-view obstructions; SF-Night (Barbarani et al., 2023), with severe illumination changes; and SVOX (Berton et al., 2021), with extreme weather and illumination variations. The results, detailed in Table 4, underscore the significant superiority of our method over previous approaches across these datasets. EffoVPR-R shows improvements of +4.3%, +0.8%, +7.9%, and +15%, +2% on Nordland, AmsterTime, SF-Occlusion, SF-Night, and SVOX-Night respectively and comparable results on SVOX-Rain. This demonstrates the high versatility of our model, which can handle extreme variations even when trained without seasonal or day-to-night changes. Figure 1 shows an examples of this case. We attribute this robustness primarily to the combination of our training method and specific re-ranking strategy over the DINOv2 model. We conduct an extensive ablation study on various hyperparameters and aspects of our approach in the Appendix.
@@
EigenPlaces 71.2 48.9 32.9 23.6 58.9 90.0
SelaVPR 85.2 55.2 35.5 38.4 89.4 94.7
CricaVPR 90.7 64.7 42.1 35.4 85.1 95.0
SALAD 76.0 58.8 51.3 46.6 95.4 98.5
BoQ 83.4 51.3 36.8 26.8 86.5 95.5
@@
5 Ablation Study
We conduct a comprehensive ablation study using two different datasets to analyze the key parameters of our EffoVPR method. Key findings are summarized below, with detailed results provided in the appendix. Analyzing the different ViT self-attention facets, we show that Value facet ($\mathcal{V}$) offers the most effective local features for VPR re-ranking (Table S3). Moreover, the selection of the layer for feature extraction is important, with the $n-1$ layer delivering the best performance, whereas using the final layer $n$ leads to a decline in results (Table S1). In Table S4 we show that our re-ranking stage remains effective even with as few as five candidates, achieving SoTA results, highlighting the method’s efficiency and effectiveness. Fine-tuning only the last five layers of the backbone network proves optimal (Figure S1). Although EffoVPR achieves state-of-the-art results with lower resolutions, increasing image resolution up to $504$px slightly improves performance (Table S7).
6 Summary
In this paper, we present both single-stage and two-stage approaches for VPR that leverage the internal layers of a foundation model for pooling and re-ranking. We challenge the necessity of recent complex architectural modifications, particularly the reliance on external pooling methods that lead to extended feature lengths in pursuit of high performance. By utilizing the model’s existing pooling layers, we demonstrate state-of-the-art results with impressive feature compactness. We introduce a novel matching method for re-ranking based on self-attention facets of a foundation model, which proves effective even in a zero-shot setting.
Demonstrating strong robustness to various appearance changes while maintaining compact features, EffoVPR offers a promising solution for tackling the VPR task in real-world, large-scale applications.